# From interpretability to inference: an estimation framework for universal approximators

### v5

@@ 4.3 Shapley regressions
$y_{i}\,=\,\Phi(x_{i})\hat{\beta}^{S}+\epsilon_{i}^{\prime}\,=\,\sum_{k\in\{0,x^{\prime}\}}\phi_{k}(\hat{f},x_{i})\,\hat{\beta}^{S}_{k}+\epsilon_{i}^{\prime}$ (18)
@@
$\mathcal{H}^{1}_{k}(\Omega)\;:\;\{\beta_{k}^{S}=1\,\big{|}\,\Omega\}\,,\,k\in\{0,x^{\prime}\}\,.$ (20)
@@ 5.1 A numerical experiment with unknown treatment interaction
$y\;=\;f_{t}(x;\alpha)+\eta\;=\;\alpha_{1}t+\alpha_{2}tz_{1}+\alpha_{3}z_{1}z_{2}+\alpha_{4}+\eta$ (21)
@@
Figure 3: True (dashed lines) versus learned (dots) treatment interaction effects $\phi_{t1}$ for RF, SVM and ANN (columns) for different sample sizes: 100 (upper row), 1000 (lower row). The best-fit linear treatment function is given for reference (solid lines).
@@ 5.2 A real-world experiment
Direct estimation is based on the specification, $y_{i}\,=\,score_{i}\,=\,\hat{f}\big{(}t_{i},z_{i};\theta\big{)}+\epsilon_{i}$. The treatment $t$ is one for subjects who saw the relatable summary, and zero for the plain text summary, with $m=1066$ ($m_{t=0}=538$, $m_{t=1}=628$). The vector $z_{i}$ contains demographic controls, especially age which we will focus in more detail below. We use cross-fitting and nested cross-validation to train and test the same off-the-shelf models (RF, SVM, ANN) as in simulation study. We also estimate a set of models following the indirect approach: causal forests (CRF; Athey and Imbens (2016); Wager and Athey (2018)), and the X-learner version of RF, SVM, and ANN from Künzel et al. (2019), where each of the prediction steps within the learner is performed by the respective base learner. We construct the treatment function (14), for which we expand model predictions to third order ($\Phi_{h}(x_{i}),\,h=3,\,x_{i}=(t_{i},z_{i})$). This guarantees that the main ($h=1$) and pairwise interactions ($h=2$) effects are the net of higher-order terms. The treatment function can be written in an abbreviated form as
$\hat{\tau}_{i}=\hat{\tau}(x_{i})=\phi_{t}+\phi_{i,t*age}+\phi_{i,t*income}+\phi_{i,t*gender}+\phi_{i,t*resid}\;,$ (24)
where terms of the form $\phi_{t*z_{k},i}$ are the estimated treatment contributions from the pairwise interaction of the treatment with the demographic characteristic $z_{k}$ of individual $i$. The term $\phi_{t*resid,i}$ is the sum of the pairwise interactions of the treatment with other controls and all other higher-order terms. We can represent the treatment function as any combination of sums of components due to the linearity of Shapley values.
@@ 5.2.1 Treatment effect estimation
We first investigate aggregate and heterogeneous treatment effects for the different models, before moving to an example of treatment interaction channels in the next subsection. The ATT obtained from all models is shown in Table 1, with the linear estimate given for reference in the last column. Looking at the upper $\hat{\tau}$-part, we see that most statistical learning models estimate the ATT well. The ANN and CRF, however, learn a somewhat subdued treatment effect. Again with the exception of these two models, the confidence intervals of all the machine learning estimates $CI_{95}^{adj}$ strongly overlap among each other and with that of the linear model, and they have comparable widths.
To asses practical aspects of our theory, we make two further comparisons at this point. First, we look at the bare treatment term $\phi_{t}$ in the treatment function (24). This should give the same ATT estimates if the treatment and control groups have the same demographic sample statistics, and if the latter (or a summary of it) has been used as the Shapley value background. This is indeed the case when comparing the upper and lower parts of Table 1. Both the estimated ATT and their confidence intervals are almost identical in almost all cases. Second, we investigate the importance of the sample split adjustment of confidence intervals for the ATT estimates. We previously noted that split variation will mostly cancel out for aggregated effects. This is the case when comparing the adjusted $CI_{95}^{adj}$ and the unadjusted $CI_{95}$ intervals in the $\hat{\tau}$-part of Table 1. This potentially allows the modeler to save on computational resources if only aggregate estimates are of interest.
@@ 5.2.2 Treatment channels: the role of age
Refer to caption Refer to caption Refer to caption
Figure 5: Estimated treatment function component $\phi_{t*age}$ for different models as a function of age for the treated based on the training bootstrap distribution: individual means (dots), best fits (solid lines, degree-4 polynomial), $95\%$ confidence intervals (CI; inner), sample split adjusted CI (outer). Three individuals older than 80 have been excluded for clearer presentation.
However, what we do observe is large variation in the distribution of estimated treatment-age interactions together with different shapes of the best-fit treatment lines, especially by based model types like the RF and (X-)SVM. We can again quantify these differences using Shapley regressions of the form (25), where we replace $\hat{\tau}_{i}$ with $\phi_{t*age,i}$. The results for this exercise are shown in Table 3 for the treated subsample.
Many models struggle to learn a relation between the treatment-age term and the outcome, since the null cannot clearly be rejected. In line with previous findings, the (X-)SVM models learn a highly significant relation. Taken together with their relatively smooth and narrow treatment functions in Figure 5 with the expected zero-intercepts, the evidence suggests that their qualitative features, e.g. the curves’ (similar) shapes, may be trusted. However, quantitative features, like estimated treatment-age values for individuals need to be treated with caution as $\hat{\beta}_{t*age}^{S}\gg 1$. These results also suggest that collecting more data, to the extent feasible, may be a fruitful way to achieve a better heterogeneous treatment effect estimation in this case.
@@ 6 Discussion
$\mathbb{E}^{B}_{test}\big{[}\hat{f}_{m}\big{]}=\frac{1}{B}\sum_{b=1}^{B}\mathbb{E}_{test}\big{[}\hat{f}_{m}^{b}\big{]}=\mathbb{E}_{test}\big{[}\hat{f}_{m}\big{]}\rightarrow f,\quad\text{as }m\rightarrow\infty\,.$ (29)
@@
$\mathbb{E}_{test}\big{[}\Phi^{b}_{m}\big{]}=\mathbb{E}_{test}\big{[}\hat{f}^{b}_{m}\big{]}=\mathbb{E}_{test}\big{[}\hat{f}_{m}\big{]}\rightarrow f=\Phi^{\star},\quad\text{as }m\rightarrow\infty\,.$ (30)
@@
Without loss of generality, let the sampling distributions of $\phi^{B}$ and $\phi^{R}$ have both $q$ elements. We order both in ascending order and form the new random variable $\phi^{joint}=\phi^{B}+\phi^{R}$ by the sum of ordered pairs. Then, at a given level $\gamma$, the confidence interval of $\phi^{joint}$ is $CI_{max}(\phi^{joint},\gamma)=[\phi^{B}_{low}+\phi^{R}_{low},\phi^{B}_{high}+\phi^{R}_{high}]$, i.e. the sum of bounds of the confidence intervals of the component variables. Since any other different relative ordering of $\phi^{B}$ and $\phi^{R}$ before forming $\phi^{joint}$ can only reduce its variance, $CI_{max}$ is the widest possible interval of $\phi^{joint}$.

### v6

@@ 4.3 Shapley regressions
$y_{i}\,=\,\Phi(x_{i})\hat{\beta}^{S}+\epsilon_{i}^{\prime}\,=\,\sum_{k\in\{0,x^{\prime}\}}\phi_{k}(\hat{f},x_{i})\,\hat{\beta}^{S}_{k}+\epsilon_{i}^{\prime}\,,$ (15)
@@
$\mathcal{H}^{1}_{k}(\Omega)\;:\;\{\beta_{k}^{S}=1\,\big{|}\,\Omega\}\,,\,k\in\{0,x^{\prime}\}\,.$
@@ 5.1 A numerical experiment with unknown treatment interaction
$y\;=\;f_{t}(x;\alpha)+\eta\;=\;\alpha_{1}t+\alpha_{2}tz_{1}+\alpha_{3}z_{1}z_{2}+\alpha_{4}+\eta\,,$ (17)
@@
Figure 3: True (dashed lines) versus learned (dots) treatment interaction effects $\phi_{t*1}$ for RF, SVM and ANN (columns) for different sample sizes: 100 (upper row), 1000 (lower row). Best-fit linear treatment functions are given by the solid lines.
@@ 5.2 A real-world experiment
[…]
@@ 5.2.1 Treatment effect estimation
We first investigate aggregate and heterogeneous treatment effects for the different models, before moving to an example of treatment interaction channels in the next subsection. The ATT obtained from all models is shown in Table 1, with the linear estimate given for reference in the last column. Looking at the upper $\hat{\tau}$-part, we see that most statistical learning models estimate the ATT well. The ANN and CRF, however, learn a somewhat subdued treatment effect. Again with the exception of these two models, the confidence intervals of all the machine learning estimates $CI_{95}^{\mathrm{adj}}$ strongly overlap among each other and with that of the linear model, and they have comparable widths.
To asses practical aspects of our theory, we make two further comparisons at this point. First, we look at the bare treatment term $\phi_{t}$ in the treatment function (20). This should give the same ATT estimates if the treatment and control groups have the same demographic sample statistics, and if the latter (or a summary of it) has been used as the Shapley value background. This is indeed the case when comparing the upper and lower parts of Table 1. Both the estimated ATT and their confidence intervals are almost identical in almost all cases. Second, we investigate the importance of the sample split adjustment of confidence intervals for the ATT estimates. We previously noted that split variation will mostly cancel out for aggregated effects. This is the case when comparing the adjusted $CI_{95}^{\mathrm{adj}}$ and the unadjusted $CI_{95}$ intervals in the $\hat{\tau}$-part of Table 1. This potentially allows the modeler to save on computational resources if only aggregate estimates are of interest.
@@ 5.2.2 Treatment channels: the role of age
Figure 5: Estimated treatment function component $\phi_{t*age}$ for different models as a function of age for the treated based on the training bootstrap distribution: individual means (dots), best fits (solid lines, degree-4 polynomials), $95\%$ confidence intervals (CI; inner), sample split adjusted CI (outer). Three individuals older than 80 have been excluded for clearer presentation.
@@
[…]
@@ 6 Discussion
$\mathbb{E}^{B}\big{[}\hat{f}_{m}\big{]}=\frac{1}{B}\sum_{b=1}^{B}\mathbb{E}\big{[}\hat{f}_{m}^{b}\big{]}=\mathbb{E}\big{[}\hat{f}_{m}\big{]}\rightarrow f,\quad\text{as }m\rightarrow\infty\,.$ (25)
@@
$\mathbb{E}\big{[}\Phi^{b}_{m}\big{]}=\mathbb{E}\big{[}\hat{f}^{b}_{m}\big{]}=\mathbb{E}\big{[}\hat{f}_{m}\big{]}\rightarrow f=\Phi^{\star},\quad\text{as }m\rightarrow\infty\,.$ (26)
@@
Without loss of generality, let the sampling distributions of $\phi^{B}$ and $\phi^{R}$ have both $q$ elements. We order both in ascending order and form the new random variable $\phi^{\mathrm{joint}}=\phi^{B}+\phi^{R}$ by the sum of ordered pairs. Then, at a given level $\gamma$, the confidence interval of $\phi^{\mathrm{joint}}$ is $CI_{max}(\phi^{\mathrm{joint}},\gamma)=[\phi^{B}_{\mathrm{low}}+\phi^{R}_{\mathrm{low}},\phi^{B}_{\mathrm{high}}+\phi^{R}_{\mathrm{high}}]$, i.e. the sum of bounds of the confidence intervals of the component variables. Since any other different relative ordering of $\phi^{B}$ and $\phi^{R}$ before forming $\phi^{\mathrm{joint}}$ can only reduce its variance, $CI_{\mathrm{max}}$ is the widest possible interval of $\phi^{\mathrm{joint}}$.